OpenStudy (anonymous):

What is the y-intercept of the equation? y = 2(3)-x

OpenStudy (anonymous):

6

OpenStudy (anonymous):

plug in a zero for x so you have y = 2*3-0 = 6

OpenStudy (anonymous):

im sorry the -x is an exponent

OpenStudy (anonymous):

thought so...:)

OpenStudy (anonymous):

procedure is still the same, plug in a zero for x: y = 2* 3^(0) = 2*1 = 2

OpenStudy (anonymous):

so what would this be then? y = 5(6.8)x 5?

OpenStudy (mayliz):

y=6-x